For this learning task, you will demonstrate your ability to solve problems using Ohm’s Law. Make sure to communicate your solutions using the GRASP approach for each question and state your final answer with the correct units and number of significant digits.


Be wary of given units like mA and kΩ – make sure to convert these to A or Ω before applying Ohm’s Law to solve the problem.


A 330 Ω resistor has a current of 11 mA flowing through it. What is the voltage drop across the resistor?

A 2.2 kΩ resistor has a voltage drop of 12 V across it. What is the current flowing through the resistor?

A resistor has a voltage drop of 8.4 V across it and a current of 150 mA flowing through it. What is the resistance of the resistor?


1
Expert's answer
2021-10-06T08:15:04-0400

The voltage drop across the 330Ω resistor with the current of 11 mA (or 0.011 A) is


"V=RI=330·0.011=3.63\text{ V}."

The current in the 2200Ω resistor is


"I=\frac{V}{R}=\frac{12}{2200}=0.005\text{ A}."

The resistance of the resistor with the current of 150 mA (or 0.15 A) is


"R=\frac VI=\frac{8.4}{0.15}=56\Omega."